An investment firm has a job opening with a salary of N45,000 for the first year. Suppose that during the next 39 years, there is a 5% raise each year. Find the total compensation over the 40-year period. 


1
Expert's answer
2021-12-19T14:15:57-0500

"s_n=\frac{a(r^n-1)}{r-1}\\

=\frac{4500(1.05^{40}-1)}{1.05-1}\\=5435989.84"